A tower is 30 feet tall. The top 25 feet part of the tower is painted orange, and bottom 14 feet of the tower is lined with stone bricks. How many feet of the tower is painted orange but is not lined with stone bricks?
Thanks!

Answer by ikleyn(52945)


First,  let us find how many feet of the tower is painted orange  AND  is lined with stone bricks.  It is

(25 + 14) - 30 = 39 - 30 = 9 feet.

Now,  let us find how many feet of the tower is painted orange  BUT  is  NOT  lined with stone bricks.  It is

25 - 9 = 16 feet.

Answer. 16 feet of the tower is painted orange  BUT  is  NOT  lined with stone bricks.

-------------------------------------------------------------------------------
Comment from student: hey why did you subtract 30 from 25 and 14 ?
-------------------------------------------------------------------------------

Good question, thank you.

Answer:  Because it is the length of the common part of the tube which is painted orange  AND  lined with stone bricks.

Make the sketch of the tube painted according to the condition,  and you will get why it is so.